Hawkins sold 10,000 shares of $1,000 par value, 6%, 5-year debenture bonds on 1/1/2017. The bonds will pay interests on 12/31 of each year. The market interest rate was 8% as of 1/1/2017. Hawkins retired these bonds for $9,580,000 on 1/1/2019.

- prepare any necessary journal entries for the followings using the effective interest method.

  1. The issuance of the bonds.
  2. Interest expense accrued in 2017 and 2018.
  3. Retirement of the bonds on 1/1/2019.
Answer #1

Answer 1:

Par value = 1000 * 10,000 = $10,000,000

Annual Interest payment = 10000000 * 6% =$600,000

Years to maturity = 5

Market interest rate = 8%

Issue amount = PV (rate,nper,pmt, fv, type)

= PV (8%, 5, -600000, -10000000, 0)

=$9,201,457.99

Discount = 10000000 - 9201457.99 = $798,541.01

Answer 2:

Year 1:

Interest expense = 9201457.99 * 8% = $736,116.64

Amortization of bond discount = 736116.64 - 600000 = $136116.64

We calculate below Year 1 and Year 2, Interest expense, Amortization of bond discount, Balance of Discount on Bonds payable and carrying value:

Journal Entries:

Answer 3:

On 1/1/2019 bonds are retired for = $9,580,000

From table (answer 2 above) carrying value of bond as on 31/12/2018 = $9484580.60

Hence:

Loss on retirement = 9580000 - 9484580.60 =$95,419.40